OpenStudy (anonymous):

Find the vertex, focus and directrix. Then draw a graph. ( x+2) squared = -6( y-1)

OpenStudy (anonymous):

this seems to be a good site http://www.scs.sk.ca/hch/harbidge/C30/Unit%201/2Parabola/2para.htm

OpenStudy (anonymous):

to get the vertex get to the standard form (you're there already). look at the bracketed terms (x+2) - vertex is at x=-2 (y-1) = vertex is at y=1 Offset for the Focus (p) is found by dividing term in front of the y. 6 in this case by a factor of 4. so offet from vertex to focus is 1.5 Because y is expessed in terms of x^2 the parabola is vertical 6(y-1) = - (x+2)^2 6y = -1 x^2 - 4x -4 + 6 therefore because the term on the x^2 term is negative the parabola opens downwards so the Focus point is (-2,1-1.5) = (-2, -0.5) So what so we know Vertex is at (-2,1) Parabola is vertical and opens downwards Focus is at (-2,-0.5) Directrix is same distance 1.5 above the vertex that the focus is below it, so it is a line where y= 2.5 See the link above for how to sketch based on the info you now have